Partitioned Matrix:

Could someone tell me why after i partitoned the matrix and do the product will get the different answer with the normal matrix product . Example partiton A and B , then AB=a, but matrix multiplication AB=c, why i will get different answer? is it normal?
Answer #1

Yes. Matrix multiplication is same for both partition and regular multiplication. While partitioning matrix it should be take care of that number of columns of first matrix should match with number of rows of second matrix that you have to multiply.

For your sake of completeness , I have solved example by two different partition. Hope , It will help you.

21 -2-6 3 1.2. AB = 1.41 (3 × 1) + (0×-3) L(-5x1) + (2x-3) (3 × 2) (0 × 4) (-5 × 2) + (2x4) 6 3 4」 -1×2-1×1 4x2 4x1 8 4 = [(-Naimol mu ltiplicaton 5 3 C-2 x1+ (6x (-3)) + (3x2 5-6+ 8 2 +18 t 6 22 25 2. t l B2 1 2.5 5 2 3 184 6 -2413 1 x 0 2921 8 22 25 22 2.5